### v1

# Data Service Maximization in Integrated Terrestrial-Non-Terrestrial 6G Networks: A Deep Reinforcement Learning Approach thanks:

Abstract
Integrating terrestrial and non-terrestrial networks has emerged as a promising paradigm to fulfill the constantly growing demand for connectivity, low transmission delay, and quality of services (QoS). This integration brings together the strengths of terrestrial and non-terrestrial networks, such as the reliability of terrestrial networks, broad coverage, and service continuity of non-terrestrial networks like low earth orbit (LEO) satellites. In this work, we study a data service maximization problem in an integrated terrestrial-non-terrestrial network (I-TNT) where the ground base stations (GBSs) and LEO satellites cooperatively serve the coexisting aerial users (AUs) and ground users (GUs). Then, by considering the spectrum scarcity, interference, and QoS requirements of the users, we jointly optimize the user association, AUE’s trajectory, and power allocation. To tackle the formulated mixed-integer non-convex problem, we disintegrate it into two subproblems: 1) user association problem and 2) trajectory and power allocation problem. Since the user association problem is a binary integer programming problem, we use the standard convex optimization method to solve it. Meanwhile, the trajectory and power allocation problem is solved by the deep deterministic policy gradient (DDPG) method to cope with the problem’s non-convexity and dynamic network environments. Then, the two subproblems are alternately solved by the proposed iterative algorithm. By comparing with the baselines in the existing literature, extensive simulations are conducted to evaluate the performance of the proposed framework.
Index Terms:
Integrated terrestrial-non-terrestrial (I-TNT) network, low earth orbit satellites (LEOSats), aerial users (AUs), Deep Deterministic Policy Gradient (DDPG).
I Introduction
With the unprecedented development of Internet of Things (IoT) technology, IoT devices and their advanced applications necessitate high throughput, comprehensive coverage, and seamless connectivity. In addition to reliable and high-speed data services, the sixth generation (6G) network is anticipated to offer end-user devices ubiquitous connectivity by integrating conventional terrestrial networks and non-terrestrial networks such as low earth orbit satellites (LEOSats). According to the recent report of SpaceX, the LEO satellites with direct-to-cell capabilities are expected to deploy more for the broader network coverage [1]. With the effective cooperation of terrestrial and satellite network operators, all users (aerial and ground users) can achieve their QoS requirements.
Despite their prominent advantages, the I-TNT networks face several challenges, such as dynamic resource management, load balancing, and interference management between terrestrial and non-terrestrial networks. There are existing works that address the interference management problem due to the spectrum sharing between terrestrial and non-terrestrial networks [2, 3, 4, 5]. For instance, the authors in [2] proposed an integrated terrestrial and satellite network where a non-orthogonal multiple access (NOMA) scheme is adopted to enhance the capacity of the ground users. Similarly, the work in [3] investigated the interference-aware sum rate maximization problem in a space-air-ground network. Considering the full frequency reuse among the LEOSats, the authors in [4] proposed a dynamic beamforming and user scheduling problem for uneven traffic demands. To alleviate the radio resource shortage problem in the I-TNT network, the authors in [5] proposed an efficient reverse spectrum sharing framework for the overall system capacity maximization. Moreover, the authors in [6] proposed a collaborated satellite and terrestrial network where the data traffic of terrestrial network is offloaded to the satellite network. By deploying the unmanned aerial vehicles (UAVs) as the relay nodes, the work in [7] proposed a learning-based resource allocation algorithm.
The aforementioned works [2, 3, 4, 5, 6, 7] only focus on interference-aware resource management for the ground users (GUs) and did not consider the existence of aerial users (AUs). Therefore, in this work, we study an I-TNT network where the ground base stations (GBSs) and LEOSats collaboratively serve the coexisted AUs and GUs in the downlink by sharing the same spectrum resources. However, due to the mobility of LEOSats and AUs, it is challenging to allocate the radio resources to all the users (both aerial and ground users) in real time. Moreover, the trajectories of AUs should be properly optimized to mitigate the severe interference in the network. It is also challenging for users to determine the association between GBSs and LEOSats to avoid high interference and maintain good network connectivity. Hence, in this work, we jointly optimize the user association, AU’s trajectories, and the power allocation of GBSs and LEOSats to maximize the data service (total downloaded data size) of the users.
II System Model
We propose an I-TNT network as depicted in Fig. 1, where the GBSs and LEOSats cooperatively serve the AUs and GUs for the downlink communication. The set of GBSs and LEOSats are defined as $\mathcal{N}=\{1,2,...,N\}$ and $\mathcal{S}=\{1,2,...,S\}$, respectively. For simplicity, the GBSs and LEOSats are denoted as base stations (BSs) defined by a set $\mathcal{B}=\mathcal{N}\cup\mathcal{S}=\{1,2,...,B\}$. In the considered area, there is a set of AUs =$\mathcal{I}=\{1,2,...,I\}$ and a set of GUs $\mathcal{J}=\{1,2,...,J\}$. Similarly, the set of all users is defined as $\mathcal{U}=\mathcal{I}\cup\mathcal{J}=\{1,2,...,U\}$, where each user can access either GBS or LEOSat for data services such as high-precision map downloading, infotainment services, or other mission-related data services. The set of subchannels that each BS can occupy in the proposed I-TNT network is defined as $\mathcal{K}=\{1,2,...,K\}$. Due to the spectrum scarcity, the GBSs and LEOSats are supposed to share the same frequency band to serve their associated users. However, the orthogonal frequency division multiple access (OFDMA) framework is implemented at each BS while serving its associated users. Therefore, there is no intra-cell interference; the users will only experience the inter-cell interference. Furthermore, the time interval is divided into a number of time slots defined by a set $\mathcal{T}=\{1,2,...,T\}$. At the beginning of each time slot, the information of users including their positions and demand is sent to the gateway with the help of GBSs and LEOSats. Then, the optimized decisions are determined at the gateway connected to the core network and informs back to the GBSs and LEOSats.
Figure 1: Integrated terrestrial-non-terrestrial 6G network.
In this work, we adopt a free-space path loss model for the communication links between AUs-GBSs/LEOSats and GUs-LEOSats as in [8] and [9]. However, the small-scale Rayleigh fading model is considered for the links between GUs and GBSs. Under these assumptions, the channel gain between AU $i$ and BS $b$ at time slot $t$ is given as
$h_{i,b}[t]=\frac{h_{0}}{(\lVert z_{i}[t]-q_{b}\rVert^{2}+(h_{i}[t]-h_{b})^{2})},$ (1)
where $z_{i}[t]=(x_{i}[t],y_{i}[t])$ and $q_{b}=(x_{b},y_{b})$ are the two-dimensional coordinates of AU $i$ and BS $b$. $h_{i}[t]$ and $h_{b}$ are the altitudes of AU $i$ and BS $b$, respectively. $h_{0}$ is the channel gain at reference distance of $1$ m. Similarly, the channel gain between GU $j$ and LEOSat $s$ at time slot $t$ can be expressed as
$h_{j,s}[t]=\frac{h_{0}}{(\lVert q_{j}-q_{s}\rVert^{2}+h_{s}^{2})}.\vspace{-0.2cm}$ (2)
However, the channel gain between GU $j$ and GBS $n$ at time slot $t$ can be calculated as
$h_{j,n}[t]=\frac{h_{0}\delta}{(\lVert q_{j}-q_{n}\rVert^{2}+h_{n}^{2})},\vspace{-0.33cm}$ (3)
where $\delta$ is the random variable that accounts for small-scale Rayleigh fading.
Since the subchannels are shared among the BSs to serve the users, the signal to interference plus nois ratio (SINR) of user $u$ when it associates with BS $b$ in time slot $t$ is given as
$\gamma_{u,b}^{k}[t]=\frac{|h_{u,b}[t]|^{2}p_{b}^{k}[t]}{I_{u,b}^{k}[t]+\sigma^{2}},\vspace{-0.2cm}$ (4)
where $I_{u,b}^{k}[t]=\sum_{\begin{subarray}{c}b^{\prime}\in\mathcal{B},\\ b^{\prime}\neq b\end{subarray}}|h_{u,b^{\prime}}[t]|^{2}p_{b^{\prime}}^{k}[t]$ is the inter-cell interference from other BSs and $p_{b}^{k}[t]$ is the transmit power of BS $b$ on subchannel $k$. $h_{u,b}[t]$ is the channel gain between user $u$ and BS $b$ and $\sigma^{2}$ is the noise. Therefore, the achievable data rate of user $u$ when it associates with BS $b$ on subchannel $k$ in time slot $t$ can be given as
$R_{u,b}^{k}[t]=w_{k}\log_{2}\big{(}1+\gamma_{u,b}^{k}[t]\big{)},\vspace{-0.25cm}$ (5)
where $w_{k}$ is the channel bandwidth.
User Association. Assuming that each user can only be assigned with one subchannel in each time slot, the association variable is defined as $\alpha_{u,b}^{k}[t]\in\{0,1\}$; its value is $1$ if user $u$ is associated with BS $b$ on subchannel $k$ in time slot $t$; 00, otherwise. Finally, the data service of the users is defined as the amount of data downloaded in time slot $t$, which is given as
$D^{\mathrm{tot}}[t]=\tau\sum_{b=1}^{B}\sum_{u=1}^{U}\sum_{k=1}^{K}\alpha_{u,b}^{k}[t]R_{u,b}^{k}[t],\vspace{-0.2cm}$ (6)
where $\tau$ is the duration of time slot $t$.
III Problem Formulation
[…]
IV Proposed Solution
This section presents our solution approach to solve the data service maximization problem in P. The formulated optimization problem in P is a mixed-integer non-convex problem because of the coupling of the binary association variables and continuous trajectory and power allocation variables. Therefore, we decompose problem P into two tractable subproblems, namely, 1) association problem and 2) trajectory and power allocation problem. Then, the two subproblems are iteratively solved to obtain the solution of P.
IV-1 Association Problem
[…]
IV-2 Trajectory and Power Allocation Problem
[…]
IV-3 State
The state space $s(t)$ of the system contains the association of users with BSs, the position of AUs, and the QoS requirements of all users in the current time slot.
1 Initialize the parameters $\phi^{\mu}$ and $\phi^{\mu^{\prime}}$ of the policy network $\mu(.)$ and target policy network $\mu^{\prime}(.)$, and set $\phi^{\mu^{\prime}}=\phi^{\mu}$;
2 Initialize the parameters $\phi^{Q}$ and $\phi^{Q^{\prime}}$ of the Q-network $Q(.)$ and target Q-network $Q^{\prime}(.)$, respectively, and set $\phi^{Q^{\prime}}=\phi^{Q}$;
3 for episode $=1$ to $E$ do
4       Initialize the random process $\Omega$ for the exploration of the action;
5       Observe state $s(1)$;
6       for $t\in\mathcal{T}$ do
7             while $C4$-$C9$ are not met and until the maximum step has reached do
8                   According to the current policy and exploration noise, select the action $a(t)=\mu\big{(}s(t)|\phi^{\mu}\big{)}+\Omega_{t}$;
9                   if The constraints are deviated then
10                         Cancel the action and apply the penalty;
11                   end if
12                  Store the transition $(s(t),a(t),r^{\prime}\big{(}s(t),a(t)\big{)},s(t+1))$ in experience replay buffer $B$;
13             end while
14            Random minibatch of $H$ transitions $(s_{i},a_{i},r_{i},s_{i+1})$ from $B$ are sampled;
15             Compute the target Q-value $y(i)=$
$r^{\prime}(i)+\gamma Q^{\prime}\bigg{(}s(i+1),\mu^{\prime}\big{(}s(i+1)|\phi^{\mu^{\prime}}\big{)}|\phi^{Q^{\prime}}\bigg{)};$
16             Update the weights $\phi^{Q}$ of $Q(.)$ by minimizing the loss function
$L(\phi^{Q})=\frac{1}{H}\sum_{i=1}^{H}\bigg{(}y(i)-Q\big{(}s(i),a(i)|\phi^{Q}\big{)}\bigg{)};$
17             Update the target Q-network parameters: $\phi^{Q^{\prime}}=\eta\phi^{Q}+(1-\eta)\phi^{Q^{\prime}}$;
18             Update the weights $\phi^{\mu}$ of the policy network $\mu(.)$ by applying the gradient method as in (16);
19             Update the target policy network: $\phi^{\mu^{\prime}}=\eta\phi^{\mu}+(1-\eta)\phi^{\mu^{\prime}}$;
20             Obtain the optimal flying velocity, directional angle, and height of the AUs;
21             Obtain the positions of AUs by $x_{i}[t+1]=x_{i}[t]+\tau v_{i}^{h}[t]\cos{(\theta_{i}[t])}$, $y_{i}[t+1]=y_{i}[t]+\tau v_{i}^{h}[t]\sin{(\theta_{i}[t])}$ ;
22       end for
23      Obtain $\bm{z}^{*}$ and $\bm{p}^{*}$.
24 end for
Algorithm 1 DDPG-Based Trajectory and Power Optimization Algorithm
IV-4 Action
The action space consists of the trajectory of AUs and the power allocation of BSs. Here, we consider the horizontal flying velocity $v_{i}^{h}[t]$, directional angle $\theta_{i}[t]$, and the height $h_{i}[t]$ of AUs as their trajectories. Therefore, the action space is defined as $a(t)=\{v_{i}^{h}[t],\theta_{i}[t],h_{i}[t],p_{b}^{k}[t],\forall i\in\mathcal{I},k\in\mathcal{K},b\in\mathcal{B}\}$.
IV-5 Reward
By taking into account the flying energy of AUs, we define the reward function for P2 as
$r(t)=D^{\mathrm{tot}}[t]-\beta\sum_{i=1}^{I}E_{i}^{\textrm{fly}}[t],$ (7)
where $\beta$ is the weight parameter, and the flying energy of AUs can be calculated by [11] and [12],
$E_{i}^{\textrm{fly}}[t]=\tau\Biggr{[}A_{0}\bigg{(}1+\frac{3(v_{i}^{h}[t])^{2}}{U_{\mathrm{tip}}^{2}}\bigg{)}+\frac{1}{2}\psi_{0}\tilde{r}\rho G(v_{i}^{h}[t])^{3}\\ +A_{1}\bigg{(}\sqrt{1+\frac{(v_{i}^{h}[t])^{4}}{4v_{o}^{4}}}-\frac{(v_{i}^{h}[t])^{2}}{2v_{0}^{2}}\bigg{)}^{\frac{1}{2}}+A_{2}v_{i}^{v}[t]\Biggr{]},$
where $A_{0}$ and $A_{1}$ are constants for blade profile power and induced power in the hovering state. $A_{2}$ is the ascending/descending power related constant and $U_{\mathrm{tip}}$ is the tip speed of rotor blade. $\psi_{0}$ and $\tilde{r}$ are the fuselage drag ratio and rotor solidity, respectively. $\rho$ and $G$ are the air density and rotor disc area, respectively. $v_{0}$ is the mean rotor-induced velocity for hovering. $v_{i}^{v}[t]$ is the vertical flying velocity of AU $i$ that can be calculated as $v_{i}^{v}[t]=\frac{|h_{i}[t]-h_{i}[t-1]|}{\tau}$.
To guarantee that all the constraints (7e)-(7j) are met, we modify the reward function in (7) as follows:
$r^{\prime}(t)=r(t)-a_{1}-a_{2}-a_{3},$ (8)
where $a_{1}$ and $a_{2}$ are penalty terms that reflect constraints (7e) and (7f). $a_{3}$ is the penalty term for trajectory constraints in (7g)-(7j). Then, the policy network generates the action value by learning the network parameter $\phi^{\mu}$ as follows:
$a(t)=\mu\big{(}s(t)|\phi^{\mu}\big{)}.$ (9)
After choosing the action based on the current state of the environment, the deep Q-network evaluates the selected action by using the Q-value as
$Q\big{(}s(t),a(t)\big{)}=\mathop{\mathbb{E}}\big{[}r^{\prime}\big{(}s(t),a(t)\big{)}+\gamma Q\big{(}s(t+1),a(t+1)\big{)}\big{]}.$ (10)
By implementing the target Q-network, the Q-network is trained to minimize the loss between them. The policy network parameter $\phi^{\mu}$ can be updated by the gradient method as
$\begin{split}\nabla_{\phi^{\mu}}J(\mu)&\approx\frac{1}{H}\sum_{i}^{H}\bigg{(}\nabla_{a}Q\big{(}s,a|\phi^{Q}\big{)}|_{s=s(i),a=\mu\big{(}s(i)\big{)}}\\ &\nabla_{\phi^{\mu}}\mu\big{(}s|\phi^{\mu}\big{)}|_{s(i)}\bigg{)},\end{split}$ (11)
where $J(\mu)$ is the output of the policy network. The proposed DDPG-based trajectory and power allocation algorithm is given in Algorithm 1.
To obtain the solution of P, we solve two subproblems iteratively. The computation complexity of our proposed data service maximization problem is $\mathcal{O}\bigg{(}\hat{r}T\big{(}UB+E\sum_{\hat{h}=1}^{\hat{H}-1}\hat{\iota}_{\hat{h}}\hat{\iota}_{\hat{h}+1}\big{)}\bigg{)}$, where $\hat{H}$ is the number of hidden layers and $\hat{\iota}_{\hat{h}}$ is the number of neurons in each layer. $\hat{r}$ is the iteration numbers.
V Numerical Results
In this section, we provide the numerical results to verify the performance of the proposed algorithm in the I-TNT network. We consider the square area of $1.3$ km $\times$ $1.3$ km, with $4$ GBSs and $2$ LEOSats. The heights of the LEOSats are set as $[300,700]$ km. The range of the values of $v_{i}^{h}[t]$, $\theta_{i}[t]$, and $h_{i}[t]$ are set as $[10,20]$ m/s, $[\frac{-5\pi}{12},\frac{5\pi}{12}]$, and $[200,300]$ m, respectively. The coordinates of GUs are randomly generated in the considered area. Moreover, the maximum transmission power of BSs is set as $[30,50]$ W and the subchannel bandwidth is set as $180$ kHz. Then, several baselines are implemented to compare with the proposed algorithm as follows:
Figure 2: Total downloaded data with different number of users.
  • Double deep Q-learning-based trajectory optimization (T-DDQN): In this baseline, the user association and power allocation are addressed by our proposed algorithm while the DDQN algorithm solves the trajectory optimization problem.
  • Random association (RA): In this framework, we randomly associate the users with BSs and only optimize the AU’s trajectory and BS’s power allocation by using our proposed DDPG-based algorithm.
  • Equal power allocation (EPA): Furthermore, in this baseline, the users are assigned with equal power values by the BSs while our proposed algorithm solves the user association and trajectory optimization.
Figure 3: Cumulative reward for different power values.
(a) Proposed.
(b) T-DDQN.
Figure 4: Optimal trajectories of AUs under different algorithms.
In Fig. 2, we illustrate the total data service of the proposed network for different number of users. To show the superior performance of our proposed algorithm, we implement several baselines such as T-DDQN, RA, and EPA. As we can see from Fig. 2, the total data service received by the users in RA is the minimum. This is because the users cannot establish a better link by the random association of BSs. On the other hand, the performance of EPA is better than that of T-DDQN. The reason is that our proposed DDPG-based trajectory optimization algorithm in EPA can determine the optimal values of two-dimensional coordinates and the altitudes of AUs to enhance the users’ data service by properly reducing network interference as well as enhancing the link quality. Nevertheless, our proposed approach can yield the maximum data service for all the users in the proposed I-TNT network.
By considering the different power values of BSs, we evaluate the cumulative reward of the proposed data service maximization problem in Fig. 4. As presented in Fig. 4, the greater cumulative reward can be obtained with the increasing power values. This is because the higher data service can be obtained with the greater power value which in turn increases the reward as given in (7). Moreover, we can observe that our proposed approach can guarantee the convergence of cumulative reward in all cases.
The optimal trajectories of AUs are illustrated in Fig. 4, where we consider $4$ AUs in the considered area. We compare our proposed approach with T-DDQN in which the continuous action space is transformed into the discrete ones. As we can observe from Fig. 4(a), the AUs in our proposed method fly near the GBSs while traveling from their starting points to the destination points. As a result, the AUs can maintain better link quality regardless of their association with BSs or LEOSats. However, as shown in Fig. 4(b), the AUs are flying near GBS-1 due to the accuracy loss of action space discretization. Therefore, it is difficult for the AUs to meet their QoS requirements. Moreover, their energy consumption will be extremely high to reach their destination points. This suggests that our proposed DDPG-based algorithm outperforms T-DDQN regarding the AUs’ trajectories.
VI Conclusion
In this paper, we have studied an I-TNT network where the GBSs and LEOSats serve the coexisted AUs and GUs for downlink communication. Considering the spectrum scarcity, interference, and QoS requirements of the users, we have formulated the problem of maximizing the total downloaded data size of the users in the network by jointly optimizing the user association, AUs’ trajectories, and the BSs’ power allocation. Then, the formulated mixed-integer non-convex optimization problem is divided into two subproblems: 1) user association and 2) trajectory and power allocation problem. The CVXPY toolkit is used to solve the first subproblem, while the DDPG-based method is employed to address the trajectory and power allocation problem. Then, the two subproblems are iteratively solved until convergence is reached. Through extensive simulation results, we have verified that our proposed framework outperforms the baseline methods such as T-DDQN, RA, and EPA.